OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? A) y = 3x + 5 B) y = 3x + 3 C) y = 3x - 17 D) y = 3x - 5

OpenStudy (australopithecus):

y = mx + b

OpenStudy (australopithecus):

where m = slope and b = y intercept

OpenStudy (australopithecus):

plug x into one of the equations which one gives you an out put of 6 or y = 6

OpenStudy (australopithecus):

remember (x,y) so you have x = 1

OpenStudy (australopithecus):

follow??

OpenStudy (anonymous):

im not following /:

OpenStudy (anonymous):

slope/gradient of 3 which means m=3 in the general equation of y=mx + c the c= y-intercept

OpenStudy (anonymous):

we are also provided with the point of (1,6) and we know m=3 so sub both of these into the general equation of y=mx + c to find c(the y-intercept)

OpenStudy (anonymous):

then we can rewrite the whole equation and easily figure out which one it is.

OpenStudy (australopithecus):

this is a graph of a line y = 3x + 5 if we sub 1 into the x value y = 3(1) + 5 y = 8 thus we have the point x , y (1,8) therefore this equation is not the one we are looking for y = mx + b is just the standard formula of a line b is just the number that occurs when x = 0 so the y intercept looks like this (0,y)

hero (hero):

y = mx + b 6 = 3(1) + b All you have to do is solve for b

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

i understand now, i dont know why i didnt before its quite simple! thanks everyone!!